# Counting rational points on smooth quartic and quintic surfaces

### v1

@@ 1. Introduction
$\rho(\tilde{C})\leq\varepsilon\log B.$
For $B\leq B_{0}(\varepsilon)$, we have $h_{F}(J)\ll_{X}\log B_{0}(\varepsilon)$. Since the number of abelian varieties of bounded height is finite, there exists some constant $c_{1}>0$ such that
$1+\rho(\tilde{C})\leq c_{1}\varepsilon\log B$
for all $\tilde{C}$, thus
$\tilde{C}(K)\leq c(g)^{1+\rho(\tilde{C})}\leq c(g)^{c_{1}\varepsilon\log B}=B^{\varepsilon c_{1}\log c(g)}.$
Since $g\leq\frac{(d-1)(d-2)}{2}$, we have finitely many possibilities for the constant $c(g)$. By adjusting the constant $B_{0}(\varepsilon)$ if needed, we obtain
$\tilde{C}(K)\ll_{X,K,\varepsilon}B^{\varepsilon},$
as we wanted. We conclude that the total contribution of these curves to the number of rational points of $X$ is $O_{X,K,\varepsilon}(B^{4/3+\varepsilon})$.
If $g(C)=1$, then either $\tilde{C}(K)=\emptyset$ or $\tilde{C}$ is an elliptic curve. In the first case, there is nothing to do. Suppose that we are in the second case. We will use the following result:
@@
For any smooth plane cubic curve $C$, we have $N_{C}(B)\ll_{K}(\log B)^{1+\rho(\tilde{C})/2}$.
Proof.
[…]
Using again the Rank Hypothesis, we conclude that $N_{C}(B)=O_{X,K,\varepsilon}(B^{\varepsilon})$ for any $\varepsilon>0$, so the total contribution of points coming from these curves is $O_{X,K,\varepsilon}(B^{4/3+\varepsilon})$.
All that is left is to bound the number of rational points in curves of genus 0 that lie inside $X$. If $C\subseteq X$ is a curve of genus 0, then either $C(K)=\emptyset$ or $C$ is a rational curve. In the first case, there is nothing to do. In the second case, let $e$ be the degree of $C$. Then $1\leq e\leq d$. If $e=1$, then $C$ is a line, and by definition it does not contribute to the calculation of $N_{X^{\prime}}(B)$. Suppose $e\geq 2$. By [paredes, Theorem 1.8],
$N_{C}(B)\ll_{X,\varepsilon}B^{2/e+\varepsilon}\leq B^{1+\varepsilon}.$
Since the contribution of a single rational curve of degree $e\geq 2$ is less than $O_{X,K,\varepsilon}(B^{4/3+\varepsilon})$, and since we have $O_{d}(1)$ possibilities for $e$, we conclude the proof of Theorem 1.3 if we show the following:
@@
Let $X\subseteq\mathbb{P}^{3}$ be a smooth surface of degree $d\geq 4$. Fix a positive integer $e$. Then $X$ contains finitely many rational plane curves of degree $e$.
@@
Consider the Hilbert Scheme $\operatorname{Hilb}_{X}^{P}$ that parameterizes the subvarieties of $X$ with Hilbert Polynomial $P$. It is a projective scheme (see [kollar, Section I.1]). Notice that the irreducible plane curves of degree $e$ inside $X$ form an open set $\operatorname{Irr}_{X}^{e}\subseteq\operatorname{Hilb}_{X}^{P}$. Let $\operatorname{Rat}_{X}^{e}\subseteq\operatorname{Irr}_{X}^{e}$ be the set of rational curves of degree $e$ inside $X$. By (2), being a rational curve is a closed condition in $\operatorname{Irr}_{X}^{e}$. Hence $\operatorname{Rat}_{X}^{e}$ is a closed subvariety of $\operatorname{Irr}_{X}^{e}$, and therefore it is a quasi-projective scheme. In particular, it has finitely many irreducible components.
@@
However, as $X$ is a smooth surface of degree $d\geq 4$, it is either a K3 surface (if $d=4$) or a surface of general type (if $d\geq 5$). In the first case, $X$ has Kodaira dimension 0, and in the second case it has Kodaira dimension $2$. This contradiction proves the lemma. ∎

### v2

@@ 1. Introduction
$1+\rho(\tilde{C})\ll_{K,d}\varepsilon(h_{K}(F)+\log B).$
Notice that, since we are assuming $H_{K}(F)\ll_{K}B^{d\binom{d+3}{3}}$, we have $h_{K}(F)\ll_{K,d}\log B$. Hence, for $B\leq B_{0}(\varepsilon)$, we have $h_{\operatorname{Falt}}(J)\ll_{K,d}\log B_{0}(\varepsilon)$. Since the number of Abelian varieties of bounded height is finite, there exist constants $c_{1}(K,d),c_{2}(K,d,\varepsilon)>0$ such that
$1+\rho(\tilde{C})\ll_{K,d}c_{1}(K,d)\varepsilon(h_{K}(F)+\log B)+c_{2}(K,d,\varepsilon)$
for all $\tilde{C}$. Thus, by the Uniform Faltings’s Theorem:
$\#\tilde{C}(K)\leq c(g)^{1+\rho(\tilde{C})} \displaystyle\leq$ $c(g)^{c_{1}(K,d)\varepsilon(h_{K}(F)+\log B)+c_{2}(K,d,\varepsilon)}$
$=$ $c(g)^{c_{2}(K,d,\varepsilon)}\cdot c(g)^{c_{1}(K,d)\varepsilon(\log(H_{K}(F)B))}$
$\ll_{K,d,\varepsilon}$ $(H_{K}(F)B)^{\varepsilon c_{1}(K,d)\log c(g)}$
$\ll_{K,d,\varepsilon}$ $B^{\big(d\binom{d+3}{3}+1\big)\varepsilon c_{1}(K,d)\log c(g)}.$
Adjusting the constant $B_{0}(\varepsilon)$ if needed, we obtain
$\tilde{C}(K)\ll_{K,d,\varepsilon}B^{\varepsilon},$
as we wanted.
If $g(\tilde{C})=1$, then either $\tilde{C}(K)=\emptyset$ or $\tilde{C}$ is an elliptic curve. In the first case, there is nothing to do. Suppose that we are in the second case. We will use [6, Proposition 1.4], which we restate here:
@@
Assume that the Rank Hypothesis holds. Let $E\subseteq\mathbb{P}^{n}$ be a smooth curve of genus $1$. Then, for any $\varepsilon>0$, we have $N_{E}(B)\ll_{K,d,m,\varepsilon}B^{\varepsilon}$.
If $C$ has no smooth points of height smaller than or equal to $B$, then $N_{C}(B)\ll_{d}1$, since $C$ has $O_{d}(1)$ singular points. Otherwise, fix a smooth point $x_{0}\in C$ with $H_{K}(x_{0})\leq B$. Then, by applying Riemann-Roch to divisors of the form $m(x_{0})$, we can find an embedding $\tilde{C}\xhookrightarrow{}\mathbb{P}^{2}$ with respect to which
$h(\tilde{C})\ll h(C)+\log H_{K}(x_{0})\ll\log B,$
where h(C~)h($\tilde${C}) and $h(C)$ denote the naive heights of $\tilde{C}$ and $C$ with respect to the embeddings $\tilde{C}\xhookrightarrow{}\mathbb{P}^{2}$ and $C\xhookrightarrow{}\mathbb{P}^{3}$, respectively.
We can take $m=2$ in the above lemma to get $N_{\tilde{C}}(B)\ll_{K,d,\varepsilon}B^{\varepsilon}$, with the height induced by the map $\tilde{C}\xhookrightarrow{}\mathbb{P}^{2}$. The above comparison of heights allows us to conclude that $N_{C}(B)=O_{K,d,\varepsilon}(B^{\varepsilon})$ for any $\varepsilon>0$. Therefore, the total contribution of the points that come from these curves is $O_{K,d,\varepsilon}(B^{4/3+\varepsilon})$.
All that is left is to bound the number of rational points in curves of genus 0 that lie inside $X$. If $C\subseteq X$ is a curve of genus 0, then either $C(K)=\emptyset$ or $C$ is a rational curve. In the first case, there is nothing to do. In the second case, let $e$ be the degree of $C$. Then $1\leq e\leq d$. If $e=1$, then $C$ is a line, and by definition it does not contribute to the calculation of $N_{X^{\prime}}(B)$. Suppose $e\geq 2$. By [15, Theorem 1.8],
$N_{C}(B)\ll_{K,d,\varepsilon}B^{2/e+\varepsilon}\leq B^{1+\varepsilon}.$
Since the contribution of a single rational curve of degree $e\geq 2$ is less than $O_{K,d,\varepsilon}(B^{4/3+\varepsilon})$, and since we have $O_{d}(1)$ possibilities for $e$, we conclude the proof of Theorem 1.3 if we show the following:
@@
Let $X\subseteq\mathbb{P}^{3}$ be a smooth surface of degree $d\geq 4$. Fix a positive integer $e$. Then $X$ contains finitely many rational plane curves of degree $e$. Furthermore, the number of such curves is $O_{d}(1)$.
@@
Consider the Hilbert Scheme $\operatorname{Hilb}_{X}^{P}$ that parameterizes the subvarieties of $X$ with Hilbert Polynomial $P$. It is a projective scheme (see [11, Section I.1]). Notice that the irreducible plane curves of degree $e$ inside $X$ form an open set $\operatorname{Irr}_{X}^{e}\subseteq\operatorname{Hilb}_{X}^{P}$. Let $\operatorname{Rat}_{X}^{e}\subseteq\operatorname{Irr}_{X}^{e}$ be the set of rational plane curves of degree $e$ inside $X$. By (2), being a rational curve is a closed condition in $\operatorname{Irr}_{X}^{e}$. Hence $\operatorname{Rat}_{X}^{e}$ is a closed subvariety of $\operatorname{Irr}_{X}^{e}$, and therefore it is a quasi-projective scheme. In particular, it has finitely many irreducible components.
@@
However, as $X$ is a smooth surface of degree $d\geq 4$, it is either a K3 surface (if $d=4$) or a surface of general type (if $d\geq 5$). In the first case, $X$ has Kodaira dimension 0, and in the second case it has Kodaira dimension $2$. This contradiction proves the finiteness of $\operatorname{Rat}_{X}^{e}$.
To conclude the proof of the Lemma, it suffices to show that the degree of $\operatorname{Rat}_{X}^{e}$ is $O_{d}(1)$. Define $\operatorname{Rat}_{\mathbb{P}^{3}}^{e}$ as the set of rational plane curves of degree $e$ inside $\mathbb{P}^{3}$. By the same argument as above, $\operatorname{Rat}_{\mathbb{P}^{3}}^{e}$ is a quasi-projective scheme. Clearly, $\operatorname{Rat}_{X}^{e}=\operatorname{Rat}_{\mathbb{P}^{3}}^{e}\cap\operatorname{Hilb}_{X}^{P}$, hence by Bézout’s Theorem it suffices to show that the degree of $\operatorname{Hilb}_{X}^{P}$ is $O_{d}(1)$.
Let $N=\binom{d+3}{3}-1$, and $\mathbb{P}^{N}$ be the moduli space of surfaces of degree $d$ contained in $\mathbb{P}^{3}$. Consider the incidence variety
$I=\{(Y,C)\in\mathbb{P}^{N}\times\operatorname{Hilb}_{\mathbb{P}^{3}}^{P}\colon C\subseteq Y\}.$
We have a projection $p\colon I\to\mathbb{P}^{N}$, with respect to which $p^{-1}(X)\cong\operatorname{Hilb}_{X}^{P}$. As the degree of a variety is stable by algebraic deformation, we see that $\operatorname{Hilb}_{X}^{P}$ has degree $O_{d}(1)$, as we wanted. ∎
@@
3. Proof of Theorem 1.4
We can adapt our approach to bound the number of rational points of bounded height on surfaces lying inside a larger projective space. Let $X\subseteq\mathbb{P}^{n}$ be a non-degenerate smooth surface of degree $d$ over $K$. The main difference here is that the ideal of $X$ is generated by several forms. Instead of Lemma 2.1, we use [2, Lemma 5], and compare $\lVert\cdot\rVert$ with $H_{K}(\cdot)$ as before by using the sets of representatives $Z_{m}$.
By [19, Theorem 2], the set of points of $X(K)$ of height smaller than $B$ can be covered by hyperplanes of height $O_{K,n}(B^{1/n})$, and by [19, Theorem 3] there are $O_{K,n}(B^{\frac{n+1}{n}})$ such planes.
Since $X$ is non-degenerate, for each such hyperplane $\Pi$ the intersection $X\cap\Pi$ is a curve $C$ of degree $d$. By Castelnuovo’s bound, $p_{a}(C)\ll_{n,d}1$, where $p_{a}(C)$ denotes the arithmetic genus of $C$. By [7, Example 3.9.2],
$g(C)=p_{a}(C)-\sum_{p}\frac{m_{p}(m_{p}-1)}{2},$
This shows that the number of singularities of $C$ is $O_{n,d}(1)$. The proof now follows identically to the proof of Theorem 1.3, and it suffices to show the following generalization of Lemma 2.3:
Lemma 3.1.
Let $X\subseteq\mathbb{P}^{n}$ be a non-uniruled surface of degree $d$. Fix a positive integer $e$. Then $X$ contains finitely many rational curves of degree $e$. Furthermore, the number of such curves is $O_{n,d}(1)$.
Proof.
It suffices to apply the argument in the proof of Lemma 2.3 to any possible Hilbert polynomial of a rational curve of degree $e$ contained in $X$. By Castelnuovo’s bound, there are finitely many such polynomials. We conclude, as in the proof of Lemma 2.3, that if $X$ contained infinitely many rational curves of degree $e$, then $X$ would be uniruled, proving the lemma. ∎